# Teleoscope: Exploring Themes in Large Document Sets By Example

### v1

@@ 2.2.1. Provenance
Our choice to focus on visualizing provenance comes from a desire to make an ML system that can be used by qualitative researchers who are not computer experts, yet still need to explain their use of ML system to stakeholders such as collaborators, supervisors, and clients (Arrieta et al., 2020) (DG1, DG2). Many topic model visualizations choose to directly visualize the underlying document clusters, that is, the result of the clustering process (Nikolenko et al., 2017; Rüdiger et al., 2022). In contrast, Teleoscope focuses on capturing and displaying the exploratory process. In the systems and visualization communities(Xu et al., 2020; Yuan et al., 2021), this is referred to as provenance tracking. For making sense of data, the importance is placed on creating a visual trace of a history of user’s interactions with the data to answer not just “What are the results?” but also “How did we get here?” (Hearst and Degler, 2013). A standard provenance system that we are all familiar with is undo/redo; a visualization of undo/redo could be as simple as a list of previous actions performed on the interface.
@@
Although the Teleoscope system does record every user action in a comprehensive history system, we hide most of that history from the user. Instead, we rely on the user creating small, comprehensible workflows that they commit to via grouping and connecting in a workflow graph. In Xu et. al’s taxonomy of topic model approaches (Xu et al., 2020), we are modeling coupled user and application state via an entity graph using semantic interactive visual analysis.
@@
Teleoscope uses concepts and design ideas from existing provenance systems, but differs by applying these ideas to qualitative understanding of corpora and topic modelling. Rather than capturing a full history of user actions, we chose to make provenance traces our primary manipulation objects. Therefore, researchers are directly creating the story of “how we got there” in a visual form that can be shared with collaborators, inspected, and updated (DG1, DG2).
2.2.2. Topic Modelling and Themes
With Teleoscope we depart from standard topic modelling because we are not interested in creating categorical topics, but instead are interested in supporting researchers in the first stage of developing rich, specific themes. The difference between categories and themes is subtle, but it is important to qualitative researchers. Rather than creating deductive categories, qualitative researchers are interested in inductive interpretation.
The difference is illustrated well with Cody, Rietz et al.’s excellent qualitative coding support tool (Rietz and Maedche, 2021). Cody addresses the problem of annotating specific sentences with qualitative codes (similar to tagging) by connecting keywords with set operations to create inclusion/exclusion rules. These rules are then applied across a corpus. This supports a style of coding that is fundamentally categorical, that is, the main action of analysis is to organize the different sentences into code categories. Although this is very useful from an organizational standpoint, the tendency is to create code hierarchies that capture more and more documents; by contrast, a good theme from a thematic analysis standpoint is very specific and involves interpretation. Teleoscope supports an interpretive approach since we are not attempting to support users in labeling documents; instead we focus on document relationships.
Many topic modelling visualizations take the approach of acting directly on topics. For example, Serendip (Alexander et al., 2014) and TopicSifter (Kim et al., 2019) both approach the problem of topic exploration and discovery by allowing users to filter and/or drill down into topics. For these tools, the topics are the primary objects of interaction, and the goal is to explore the topics themselves. Both are a top-down approach. Teleoscope is most conceptually similar to Scholastic by Hong et al. (Hong et al., 2022) because we also take a human-centred approach and provide a hierarchical clustering algorithm as a “machine-in-the-loop” approach. Their elegant approach of visualizing the document space differs greatly from our choice of visualizing the process of discovery. Also, as with Cody, they focus on document coding, which we do not.
Teleoscope differs from topic modelling tools by focusing on (1) discovering relevant documents rather than making claims about a whole corpus; (2) supporting collaboration, reproducibility, and schema expression; (3) co-developing themes with the computer through creative data exploration; and (4) allowing users to interact in real time with a larger corpus than other tools (DG1, DG2, DG3).
2.3. NLP Approach
@@
To facilitate the concept of themes over topics, we take the approach of modelling by example to allow users to explore and then structure the document space. Lissanddri et al. (Lissandrini et al., 2019) describe example-based search approaches as having a resurgence in popularity. A variety of example-based query systems have been introduced that attempt to synthesize the query from examples (Martins, 2019; Fariha and Meliou, 2019; Lissandrini et al., 2019). Teleoscope follows these works by using semantic similarity to drive exploration (see Fig. 4 and Section 3.1 for details). Teleoscope differs from these works by not attempting to construct deterministic queries or models of the entire corpus. Instead it relies on the interaction process and the user’s own sense of saturation to determine the extent of the exploration.
@@
When a Teleoscope user has finished a phase of exploring via semantic similarity, they can switch to structuring the document space via semi-supervised dimensionality reduction. Variations on this approach are used in recent human-centred ML and visualization systems (El-Assady et al., 2019; Meinecke et al., 2021; Asudani et al., 2023; Sperrle et al., 2021). The premise is to take a large language model and reduce dimensions along which a similarity metric is defined. Teleoscope uses the Universal Sentence Encoder (USE) for the base exogenous model, which encodes all documents as 512-vectors (Cer et al., 2018). In Teleoscope, we use grouped documents as control inputs to define the similarity metric (i.e., this is the “supervision” part of our semi-supervised topic model approach). We use Universal Manifold Approximation and Projection (UMAP) for reduction to five dimensions (McInnes et al., 2018). Clustering uses Hierarchical Density-Based clustering (HDBSCAN*) (McInnes et al., 2017).
@@
3. Teleoscope Design
Figure 2. Above are the different types of nodes and operations in the workflow graph. Document sources can be a single document, a group of documents, the results of a keyword search, or all documents in the corpus. Controls for the target Teleoscope ranking operation or the Projection clustering operation can be any subset of the corpus or a note (which includes arbitrary text). Operations output a list of document lists which are partitioned per document source.
Figure 2. Above are the different types of nodes and operations in the workflow graph. Document sources can be a single document, a group of documents, the results of a keyword search, or all documents in the corpus. Controls for the target Teleoscope ranking operation or the Projection clustering operation can be any subset of the corpus or a note (which includes arbitrary text). Operations output a list of document lists which are partitioned per document source.
Figure 3. Users start a new workflow using a keyword search. Then, they can explore documents by dragging them onto the workspace, skimming titles and reading. When they find documents they like, they can annotate and group the documents; they can further combine sources with set operations. Any source can be ranked according to controls. Teleoscope can also cluster documents for them into groups.
Figure 4. ML similarity scores are based on vectorizing text and measuring the distance between vectors. In the above example, a model would assess doc1 and doc2 to be most similar to each other because they share the phrase lazy dog. A sophisticated model, with enough training data, may be able to establish semantic similarity between sun and summer.
The similarity score between doc2 and doc3 would still be correctly measured as lower than doc1 and doc2. However, as a qualitative researcher who is performing interpretive research, you would want the ML model match your personal interpretive schema. In this example, you would want the model to emphasize the similarity between sun and summer.
Figure 5. Teleoscope encodes all documents using the same model. We chose to use the Universal Sentence Encoder (USE) model, but any model that takes text and outputs vectors could be substituted. Teleoscopes are oriented by creating a Teleoscope vector, which is an average over all vectors from nodes connected to the control input, which may include any source, i.e., documents, groups, searches, or notes. Each document in sources is ranked by distance to the Teleoscope vector. If no sources are indicated, all documents in the database are ranked. For frontend display speed, we limit the number of stored ranks.
In this section, we outline our design decisions for Teleoscope. Our design process took over two years and involved many iterations. As we used Teleoscope for our own data exploration, demonstrated Teleoscope to others, and performed initial pilot studies, our engineering and UX requirements expanded.
Our design takes the most direct inspiration from applying processes used in thematic analysis to data collection. Importantly, Teleoscope does not facilitate thematic analysis, but instead applies ideas from thematic analysis to data collection to increase theoretical and practical rigour. For thematic analysis, Braun and Clarke emphasize that during the data familiarization phase, there needs to be active engagement with the data:
{quoting}
Use whatever format works for you (e.g., annotating transcripts, writing comments in a notebook or electronic file, underling portions of data) to highlight items potentially of interest. Note-making helps you start to read the data as data. Reading data as data means not simply absorbing the surface meaning of the words on the page, as you might read a novel or magazine, but reading the words actively, analytically, and critically, and starting to think about what the data mean. (Braun and Clarke, 2012)
We contend that active engagement via annotation, direct manipulation, and exploration is vitally important to exploring an opinionated data collection at scale. Arranging and re-arranging data allows patterns to emerge; in Teleoscope, data workflows can be arranged and rearranged as well as individual documents. Further phases of thematic analysis are also reflected in the Teleoscope workflow, which are referenced below.
Creative, engaged interaction with data required systems engineering choices that made real-time feedback possible, motivating our development of a drag-n-drop interface with a distributed backend. We initially designed Teleoscope as a simple dashboard interface for manipulating keywords with a bag-of-words model, however, found that manipulation was an important part of sense-making.
After about a year of our own design iterations, we moved into a second phase of user-driven design by recruiting social scientists for an in-depth usability study, analyzed the study results with a visualization research group, and then performed a field deployment with a dedicated qualitative research group. We also designed the system to be robust enough to be used by the general public and released Teleoscope openly on the internet. Our studies and data collection are detailed in the next section. This section describes the outcome of the design iterations so that the reader is able to understand the current Teleoscope system design, which is the third major redesign and includes design up to and including the public release.
3.1. Teleoscope interface concepts
Teleoscope consists of an infinite whiteboard workspace where users connect data sources to computational operation targets to create a computational graph. There is also a sidebar for quick navigation and reading. A chain of sources are used to control the ML operations and form the basis of our workflows. From an interaction perspective, the workflows become curated digital artifacts that leave a trace of the researcher’s exploration of the data. From a computation perspective, the workflows are user-curated provenance graphs.
3.1.1. Workspace
We developed the Teleoscope workspace to reflect the process of arranging data on a table, due to conversations with qualitative researchers about needing to arrange data to make sense of it (DG2). We wanted Teleoscope to reflect a researcher’s personal meaning of their data arrangements with computational meaning in the ML model (DG1). Therefore, the Teleoscope workspace is a drag-n-drop visual workflow editor which produces computational graphs. We have two types of nodes in the workflow graph: document sources and target operations. We chose a direct manipulation metaphor for positive transfer from whiteboard apps that qualitative researchers are familiar with (such as Miro) and to mimic the process of arranging and rearranging documents on a physical desktop such as in affinity diagramming. Based on user feedback, we settled on a whiteboard workspace metaphor with a visual workflow editor. We also implemented a quick viewer in the right hand sidebar and keyboard navigation for skimming document titles and body text. See Fig. 6 for an annotated depiction of the following sections.
3.1.2. Sources
A document is the primary interaction object in Teleoscope. Documents display a title, text and metadata information within the node or in the quick viewer when selected. Best performance for documents is 1-3 paragraphs, such as social media posts. Users can review documents to decide whether they are topically relevant by grouping, bookmarking, or using them as input to operations. A search is a fuzzy keyword search across the full document set; queries are in plain text and support set operations like and, or, and not. Keyword search was provided since this is a common search interaction style that qualitative researchers have often been explicitly trained in and are used to thinking about. Documents can be sorted into labelled groups, either by dragging and dropping into a group node window, or by selecting the group in a drop down. A note can contain arbitrary text, which is then vectorized live as the user types. Users can use notes to create annotations, or they can use the note to drive teleoscope operations.
Computationally, all sources are lists of one or more vectors. Interactionally, sources are meant to be treated like (1) document collections; and (2) examples that will drive the ML system to find relevant documents according to what the user arranges to be together.
3.1.3. Targets
Targets are the ML operations placed at the end of a workflow and produce collections of document collections as outputs. Each source will create a new document collection, whereas each control will manipulate distance metrics and similarity scores. Targets are meant to reflect the user’s mental model as they explore the data so that the user can update their mental model along with the machine model.
The teleoscope operation ranks source documents relative to the average vector of all control documents (see Fig. 4). If no sources are connected to a teleoscope, it will rank all documents in the corpus; otherwise it will rank each source subset independently.
The Projection operation runs semi-supervised dimensionality reduction and clustering.
Control inputs to a projection define a distance metric: if two documents are in the same group, their distance is set to a minimum; otherwise, their distance is the cosine distance between their vector encodings. Users can choose to include an additional rule in the distance metric that sets the distance between groups to a maximum. Doing so forces the system to separate groups when clustering; otherwise, groups may be clustered together.
In the projection operation, source inputs can be used to define the domain of the documents being clustered. If no source inputs are provided, users can decide between a random subset of the corpus, or a selection of documents ranked relative to the average vector of control inputs.
3.1.4. Source and target operation: Set Operations
Set operations are provided so that users may non-destructively combine sources and inspect results. They include standard set operations of union, difference, intersection and exclusion (the complement of intersection). They are the only operations which are both sources and targets; they can be combined to create chains of operations (see Figure 6).
3.2. Workflow Design Patterns
3.2.1. Rank
The Teleoscope ranking operation can be used to explore documents from the whole corpus, but if a source is connected, the document list can be ordered relative to the control. The controls can be any source; in Fig 6, we can see a search source being ranked by a note control.
Figure 6. The above is an annotated illustration of the Teleoscope workspace, explaining workspace operations and giving example workflow patterns. Annotations are in serif fonts whereas the workspace items are sans-serif. See Figs.  1 and  3 for detailed window depictions and conceptual overviews of operations.
3.2.2. Chaining Set Operations
Set operations can be used to join and filter documents as well as intermediary connective operations. In Fig 6, we can see that set operations are chained to get a specific document source which is then ranked by the Teleoscope operation. Set operations can be used for non-destructive editing of groups, construct larger sources out of smaller sources, or to narrow down large groups. When combined with machine-generated groups after clustering, they can also give insight to the way in which the machine clusters overlap with the user-defined groups, either providing confidence in a user’s group choice or indicating an area for further exploration.
3.3. Technical systems design
Figure 7. We designed a strict dataflow for Teleoscope to aid with concurrency and to mask remote processing latency. The NextJS client uses a polling strategy with the Stale-While-Revalidate (SWR) library to pick up state changes from MongoDB. The client sends API requests to RabbitMQ and concurrent Celery workers read from RabbitMQ and write to MongoDB. Dispatch and task modules are split and can run on different virtual machines for load balancing.
Reflecting our goal of real-time interaction (DG3), Teleoscope’s backend is engineered to support continuous iteration and interaction. This was a non-trivial task which is not always supported by other systems due to the high computational workload. Through internal testing, we found a distributed backend was necessary to run Teleoscope computations without blocking user interaction. We chose to precompute, cache, and distribute as much as possible to make on-the-fly calculations seem quick. For example, a Teleoscope rank of cached data is nearly instantaneous despite sorting hundreds of thousands of documents (average 1601 ms for update on UI including network latency). For operations that are not instant, the interface allows continuous interaction with other elements. We believe that this user experience goal is vitally important to Teleoscope being capable of facilitating the flow state needed for sense-making.
From a community perspective, we wanted Teleoscope to be available for qualitative researchers who are non-computer experts, which necessitated a robust enough system to survive a production-level environment on the open internet, including almost-one-click deployment, user accounts, security, and backup systems. Teleoscope is continuously available live at [anonymized for review after consultation with CHI chairs] for use by the general public. We have an active user community on Discord where we take bug reports and design requests.
3.3.1. Frontend
Our frontend is built using the NextJS ecosystem to manage React development and deployment, ReactFlow for the graph drag-n-drop workflow system, and Material UI for the design elements, as well as a variety of smaller libraries and custom components. NextJS was chosen because of its large user community and full stack support, including data fetching and user authorization libraries. We chose ReactFlow after experimenting with a number of whiteboarding, windowing, and drag-n-drop libraries; it is also a mature and actively maintained freemium product. Material UI implements Google’s material design in React.
3.3.2. Backend
Teleoscope uses a distributed backend with RabbitMQ for messaging and Celery to execute tasks. Our database is MongoDB. To ensure continual service, our system is daemonized with native linux systemctl as well as the pm2 library for node and python applications. Both perform process monitoring and memory management.
3.3.3. Dataflow and History management
React uses a virtual document-object model (DOM) to ensure a strict dataflow model for user actions and system state. Similarly, we designed a dataflow policy such that the frontend (almost) entirely makes requests to the backend to manage system state on the server. This means that actions that mutate database state are strictly sent via a secure websocket connection to RabbitMQ through a well-defined API. With the exception of user registration, there is no direct database mutation by the frontend. Similarly, backend state updates are managed by the Stale-While-Revalidate (SWR) data caching and fetching system in NextJS.
Keeping this strict policy has benefits and drawbacks, mostly having to do with interaction availability. Any large-scale calculations and mutations are offloaded to the backend while the frontend waits for data to be marked as stale to refresh the local client cache. The trade-off is that some state changes that require a backend response may be impacted by network latency.
4. Design Studies

### v2

@@ 2.2. Teleoscope’s Visualization Approach
Our focus on visualizing provenance comes from a desire to make an ML system that can be used by qualitative researchers. These people are not computer experts, yet still need to explain their use of an ML system to stakeholders, such as collaborators, supervisors, and clients (Arrieta et al., 2020) (DG1, DG2). Many topic model visualizations choose to directly visualize the underlying document clusters, that is, the result of the clustering process (Nikolenko et al., 2017; Rüdiger et al., 2022). In contrast, Teleoscope focuses on capturing and displaying the exploratory process. In the systems and visualization communities (Xu et al., 2020; Yuan et al., 2021), this is referred to as provenance tracking. For making sense of data, the importance is placed on creating a visual trace of a history of user’s interactions with the data to answer not just “What are the results?” but also “How did we get here?” (Hearst and Degler, 2013). A standard provenance system that we are all familiar with is undo/redo; a visualization of undo/redo could be as simple as a list of previous actions performed on the interface.
@@
Although our Teleoscope system does record every user action in a comprehensive history system, we hide most of that history from the user. Instead, we rely on the user creating small, comprehensible workflows that they commit to via grouping and connecting in a workflow graph. In Xu et. al’s taxonomy of topic model approaches (Xu et al., 2020), we are modeling coupled user and application state via an entity graph using semantic interactive visual analysis.
@@
Teleoscope applies ideas from provenance research to qualitative understanding of corpora. We chose to make provenance traces our primary manipulation objects. Therefore, researchers are directly creating the story of “how we got there” in a visual form that can be shared with collaborators, inspected, and updated (DG1, DG2).
2.3. Teleoscope’s NLP Approach
@@
To facilitate the concept of themes over topics, we take the approach of modelling by example to allow users to explore and then structure the document space. Lissanddri et al. (Lissandrini et al., 2019) describe example-based search approaches as having a resurgence in popularity. A variety of example-based query systems have been introduced that attempt to synthesize the query from examples (Martins, 2019; Fariha and Meliou, 2019; Lissandrini et al., 2019). Teleoscope follows these works by using semantic vector similarity to drive exploration. Teleoscope differs from these works by not attempting to construct deterministic queries or models of the entire corpus. Instead it relies on the interaction process and the user’s own sense of information power to determine the extent of the exploration.
@@
When a Teleoscope user has finished a phase of exploring via semantic similarity, they can switch to structuring the document space via semi-supervised dimensionality reduction. Variations on this approach are used in recent human-centred ML and visualization systems (El-Assady et al., 2019; Meinecke et al., 2021; Asudani et al., 2023; Sperrle et al., 2021). The premise is to take a large language model and reduce dimensions along which a similarity metric is defined. Teleoscope can use any model for the base exogenous model and is currently using FlagEmbedding’s BGEM3 model from HuggingFace (Chen et al., 2023; Wolf, 2019). In Teleoscope, we use grouped documents as control inputs to define the similarity metric (i.e., this is the “supervision” part of our semi-supervised topic model approach). We use Universal Manifold Approximation and Projection (UMAP) for reduction to five dimensions (McInnes et al., 2018). Teleoscope clustering uses Hierarchical Density-Based clustering (HDBSCAN*) (McInnes et al., 2017).
@@
2.4. Teleoscope Workspace
The Teleoscope workspace (see Figure 2) is an infinite canvas mindmap with node-based dataflow functionality. The basic interaction premise is that data sources flow into targets. Data sources are arbitrarily-long document lists that subset the full document space. The search operation filters the full dataset based on a keyword search (with basic boolean text search operations). Groups are document lists that users have explicitly selected documents to be included in.
All document sources can also be used to control operations. If a user makes a group of documents, they can then use it to control a Rank operation by connecting the Group to the Rank target node. The Rank operation will then perform a vector similarity search of the full database based on the average of the group vectors, and return an ordered list of the most similar documents up to a threshold. If a specific source (like a keyword Search) is connected, then the Rank will only order that source. Custom annotations called Notes can also be used to control Rank operations. Notes are arbitrary text that the user enters which are immediately vectorized.
Projections are automatic clustering operations. The Projection will take a source and create a lower-dimension projection of the source documents based on the controls by making a custom distance metric. For each control, the Projection will say that there is a minimum distance between each document within the control group, and a maximum distance between documents that are in different control groups. The Projection then creates a dimensionally-reduced model of the source documents based on this custom metric, which are then clustered. The practical upshot is that the Projection operation attempts to group all source documents as if they were in the same groups that the user has defined.
Next, we take a step back and consider the underlying metaphors that Teleoscope implements. We then consider how these metaphors are enacted by a user of the tool. Throughout, we provide examples of the patterns that emerged through our usage of Teleoscope, as well as the usage of participants in our case studies (detailed in Section 4).
3. Teleoscope Metaphors and Usage Patterns
Figure 3. Large corpora in the thousands to millions of documents are difficult to make sense of, but LLMs are making it technically feasible to try. How do you find which documents are important to you in a rigorous, repeatable, sharable manner?
The design of Teleoscope was largely motivated by our own need to explore large amounts of data in a satisfying, repeatable way. In particular, before we designed Teleoscope, we were interested in exploring privacy/security concerns within Reddit’s Am I the Asshole (AITA) advice forum (Reddit.com, 2024). The AITA forum has a format in which an original poster (OP) will ask a question about a situation in which they are unsure whether they acted like an “asshole,” describe the situation, then ask the AITA community to vote on whether or not they were an “asshole” in that situation. The AITA community will discuss in large comment threads while pronouncing concretely either “You’re the Asshole” (YTA), “Not the Asshole” NTA, or the middle ground of “Everyone sucks here” (ESH). These votes are automatically tallied and included as tagged flair on the original post. Occasionally, the OP will provide further information, comment, apologize, or argue with the AITA community.
We chose this dataset because it provides many interesting, explicitly normative accounts of conflicts. The corpus is large, with nearly 1 million documents, which we filtered down to  350K usable documents by removing posts that were removed by moderators or were otherwise corrupted. Our original exploration topic was intentionally broad (Chen et al., 2018a), looking simply for general privacy/security concerns that involved technology. Eventually, we settled on “Nuanced challenges to our current privacy model” as the first large-scale theme that we would explore. This became the nucleus of our schema crystallization (defined in the next subsection): trying to decide what we actually meant by that theme. [^1] At the moment, Teleoscope does not support nested document display such as with Reddit’s comment threads.
3.1. Schema Crystallization: Starting with nucleation
Schemas are a concept in psychology that refers to a set of beliefs, attitudes, and emotional responses related to an archetypal situation, e.g., “driving.” An unstated belief might be that “all drivers are not paying attention” which is related to a dismissive attitude towards drivers, and an emotion of frustration. If a driver in front were to take a few seconds too long before moving at a green light, the holder of the schema might externalize their schema with an explanatory statement by saying out loud that the driver “must be on their phone”—whether or not that is true. After saying it out loud, a passenger may reflect and amend the externalized schema by saying “maybe they were just waiting for a car to finish turning.”
Berret and Munzner introduce the concept of schemas as part of the sense-making process with data visualization (Berret and Munzner, 2022b). Similar to the psychological concept of a schema, data visualization provides a method by which a researcher can externalize their schemas, allowing them to inspect, amend, and develop their internal schemas along with the externalized artifacts of the schema.
Similarly, crystallization is a term introduced by Ellingson (Ellingson, 2009) to extend post-positivist ideals of triangulation to post-modern qualitative research. The idea is that if you triangulate many times, you have a more complex crystalline structure. This structure gives you many facets or views on the data. We also liked the metaphorical similarity to graph visualizations that we thought about while designing Teleoscope’s data structures. Extending Ellingson’s metaphor into sense-making with very large datasets, we imagine schema crystallization as a process that happens both in a person’s mind, and with the computational representations (see Figure 3).
We define schema nucleation to introduce the metaphor of a growing crystal for schema crystallization. A large crystal grows from a small nucleus. Similarly, expressing and developing a schema must start from a small entrance into the dataset. With Teleoscope, this can happen with a standard fuzzy keyword search (see Figure 4), or by simply skimming through the dataset.
3.2. Schema Nucleation: Keyword Searches
Figure 4. During nucleation, ideas about the corpus are just starting to unfold and develop. Quick interaction is key to keeping an open mind while exploring. Crucially, each interaction is both expanding possibilities about the corpus through discovering new ideas, and potentially prematurely creating closure and opening up the possibility of confirmation bias. Keeping a visual trace of explored avenues allows for the necessary systematic challenge to one’s biases. This image includes screenshots and descriptions of our real process of nucleation. For each nucleation, we started from a vague conceptual keyword such as “privacy and discovered more concrete keywords such as “passcode.”
As our deployment case studies eventually revealed, many qualitative researchers are formally trained in methodologies that manage keyword searches, including boolean search (Namey et al., 2008). A fuzzy keyword search produces a definite set of documents (see Figure 5). Every keyword search subsets, facets, or creates a view of the data. Every skimmed document is starting to shape the researcher’s early theories of what the dataset does and does not contain.
Figure 5. Ideas start off ambiguous. As we wonder about our corpus, hunches, notions, and predictions emerge that we can test against the corpus. Our pre-existing schemas need to be developed by externalization (e.g., writing on a whiteboard). For this interface, externalization starts with guessing keywords to discover relevant documents.
We began our exploration of the AITA dataset with keyword searches for common privacy/security and technology related concepts, such as privacy, security, password, and account. Immediately, questions of biasing our results through the choice of keywords arose. How did we know that we had the right keywords? How do we determine whether the posts captured by those keywords are appropriate?
After skimming titles and reading through many documents, we started to get a general sense of curiosity about certain themes. For example, privacy brings up potentially relevant documents, such as those about changing passwords, or reading through a partner’s messages. It also brings up less relevant documents, such as using the toilet in private.
Documents that are interesting can be annotated or dragged onto the Teleoscope interface for later. We started to create groups of documents that captured vague ideas about undetermined themes. Importantly, although every interface action helped to develop our internal schemas, there was not yet any strong externalized commitment to the schema. This was a generative, playful, exploratory stage, rather than a stage with highly directed action.
With enough exploration, the interface’s workspace became a visual artifact, or concrete external representation of the ambiguous internal schema. The workspace also can become messy, with staged and annotated items in no clearly meaningful arrangement. To continue to develop the schema, a perspective needs to be imposed on the dataset, which grows a facet of the crystal.
To start to make sense of our keyword searches, we needed put them into some kind of order. With LLM embeddings, vector similarity search is an approximation of semantic similarity. However, there is a difference between one’s schematic sense of thematic similarity and the sense of similarity in a vector similarity search.
Figure 6. When developing a theme, some documents may be more or less illustrative of that theme and therefore more or less relevant to analysis for that theme. When organizing themes, documents must eventually be included or excluded from analysis, however, getting a sense for which documents are in or out takes iteration.
Machine similarity is a statistical similarity, and there is no inherent discrimination between underlying word senses, interpretations, theoretical perspectives, or meaningful poetic connections. For example, in Figure 6, the Rank operation is controlled by a document that is complaining about a mom “snooping.” There is a semantic similarity that is captured with the subsequent documents referencing “hiding” but it does not capture the underlying sense of motive or discomfort that is implied in the document—these are our interpretations which are not necessarily present in the text, but instead in our own schemas. Imposing an order based on similarity ranking is a helpful first step, but we could not assume that the machine’s proposed order would be the same as a human’s thematic grouping.
As we skimmed documents, we added new controls to better approximate the yet-undetermined themes. Rankings were updated to reflect the controls. We tested our “gut feelings” as we externalized by arranging and grouping documents before themes could be formalized in a concrete way. Eventually, we wanted to create a group of documents that were confirmed to be within a particular theme.
3.3. Determining Information Power (saturation)
Figure 7. As documents are explored, both a conceptual relevance and calculated metric relevance are revealed. Both our internal sense of thematic relevance that comes from our internal schemas and the externalized vector similarity ranking are impacted through interactions. In this diagram, both senses of relevance are depicted as pink crystal surfaces, where height corresponds to relevance.
Figure 8. Here is an example of a theme that has been faceted to the point of saturation. Many different facets are presented on the same curated group that subsets the larger corpus.
Richness and saturation are two values that are used to talk about rigour during theme development (Braun and Clarke, 2012; LaDonna et al., 2021). Qualitative researchers are also using the term “information power” to liken the value of well-triangulated data to statistical power. I.e., in statistical modelling, we value many data points, but in qualitative research we also value differing data points. Whether something is rich, saturated, or has high information power [^2] Here, we will use saturation rather than information power since the term reads as a quality rather than an amount. is defendable through rhetoric, but not clearly measurable, which makes it difficult to mechanize within a computational system. Both terms refer to having dense meaning, richness through making connections to theory, between data points, and drawing in other sources; saturation through each new data point not providing a new perspective on the theme. Teleoscope attempts to impose a computable framework on these concepts.
While thematically exploring large corpora, it is important to determine saturation because we want our themes to be rich. Remember, we are not doing a final thematic analysis with this method, rather we are trying to find the right dataset that will be able to draw out certain themes of interest. Therefore, we needed to find a way to define saturation in terms of using a system like Teleoscope.
Our approach is to liken saturation to a high level of crystallization (Tobin and Begley, 2004). Many facets are available to the researcher: Search and Rank operations have been iterated on many times, Groups have been annotated, broken into different thematic subsets, and new unskimmed or read documents are not often appearing. In this section we will explore what it means to determine saturation under the conditions of uncertainty that are intrinsic to large corpora.
3.4. Crystal facet density is comparable to Saturation
With large corpora, it is impossible to read every relevant document. Some qualitative researchers will hire large teams of coders to achieve results that are not dissimilar to a mechanized result, because the enormity of the operation requires a systematic approach. Coders are given detailed instruction manuals and have team meetings to ensure inter-rater reliability. However, it is more in line with the values of thematic analysis to have individual interpretations that are shared and develop through collaborative discussion.
In our process of exploration, we found interesting example documents that became our nucleation points. Often, they were semantically similar to other documents that we found much less interesting, but not meaningfully similar. This was a surprising distinction.
Through our initial exploration, we found large groups of documents that we found to be too similar to analyze individually. For example, searching for wifi as a keyword produces many documents that are about fairness in paying for wifi, and the social difficulties of determining who has access to wifi. After exhaustively looking through every wifi document, it was clear that documents that included the wifi keyword were thematically equivalent. There was enough to delve into a deeper analysis, but further faceting within the wifi keyword search was not needed.
However, the thematic curation was not yet done. Wifi is only one keyword. Through a Rank operation, we were able to find similar themes of family sharing of accounts such as Netflix, Spotify, and other streaming services. We made a list of keywords of unique streaming services with similar problems, similar to the qualitative researchers in our deployment case studies. Like them, we reached a point of keyword saturation. Assembling all keywords gave a list of 13,717 documents, which is too many to read. The theme had not yet fully crystallized, since there were not yet enough facets created (see Figure 8 for the fully faceted worked example).
Human-created groups indicate that the person feels that certain documents potentially belong together in a theme. When a few groups are created, a Projection operation can be effectively used to have machine-proposed groups (called “clusters”) based on the human-made groups. This provides more facets through collaborating with the machine. Clusters can be dragged onto the workspace just like groups, where they can be refined, added to, and used for other operations such as Rank to find similar documents that were not in the original Projection source.
3.5. Incorporating Annotations and Arrangements
Annotation is an important part of externalizing schemas and developing themes. Arranging documents on an interface develops an embodied thematic understanding. Just like with affinity diagrams, arranging and re-arranging allows for quick, iterative mini-themes to emerge and disappear. Eventually, themes congeal and need to be annotated to be properly understood.
Figure 9. We discovered that we could us vectorized annotations to creatively summarize and search. We got a sense for archetypal phrases that Reddit AITA posters use, coming up with example quotes that were fictitious, but typified persona. We tried this because we could not find a good combination of keyword and Rank operations on posts because of a theme we were trying to crystallize about the reliance of families (and particularly older family members) on younger family members. We could imagine what a younger family member might complain about on AITA. Along with our own generated annotations, we used quotes from documents in the dataset for semantic mixing. This helped us discover real examples as well as disconfirm our belief in the prevalence of this theme in the dataset.
Since Teleoscope allows annotations to be vectorized, this affords an interesting new mechanism of searching, which we will call searching by archetype (see Figure 9). As themes are annotated, common narratives emerge archetypally representing a theme. For example, “mom always snoops through my messages” might be a phrase that we would assign to a persona developed from posts about password sharing on mobile devices. Since they are also vectorized by the Note operation, they can be used as part of similarity searches. This is a new and interesting way of curating results. If a researcher can accurately take on the voice of an archetypal poster, they can find results that are thematically similar. Guessing and testing can help to refine the archetypal search results. Multiple archetypal phrases can be used. This is a novel search mechanism that is based in a creative arts practice that has essentially not be enabled with prior technology.
Annotations can also be quotations. As example posts start to emerge, refining the vector similarity search using quotations allows for more on-theme results. This can be creatively combined with archetypal searches as well as document similarity search for further semantic mixing.
3.6. Signposts are differentiating examples
The more we worked with our documents, the more we found that we had (1) documents sets to analyze in large batches; (2) documents sets to confidently ignore; and (3) special example posts that were “exceptions that proved the rule,” warranting special focus and attention—signposts. Signposts were closest to our internal schemas, but difficult to express as external workflows. They became key documents to determine saturation and major targets for search. By finding these hidden gems, we anchored our crystallization efforts.
Figure 10. After reading through many posts, we wanted to figure out how to differentiate our research goals from common framing of spouse surveillance to something more nuanced. This post became a differentiating example—a signpost—to generate a group of posts that we found had a level of nuance. However, it was difficult to represent this in terms of vector similarity search, and it was difficult to articulate in words to ourselves what we were interpreting in the data.
An example is “AITA for secretly muting my wife’s emails while on vacation?” This post (see Figure 10), along with other signposts, helped us to solidify our thematic curation around “ambiguous privacy/security violations that challenge the accepted value systems of current privacy models.” However, this was an extremely difficult theme to look for, only discoverable through carefully reading most of a Rank operation. That is because the content of the post is a nuanced problem about searching through a partner’s email, which rarely surfaces posts about nuanced situations.
There is no nuance detector built into vector search. Despite searching using keywords, Rank operations, and archetypal searches, there was no automatic way to bring up posts that were similar in underlying nature. This illustrates the fundamental difference between what a machine can easily do with a vector search vs. the value of a human interpreter.
Our interpretation of the post is that there are conflicting value systems at play. On the one hand, the husband is clearly violating a privacy norm and controlling his wife’s actions. On the other hand, he believes that the is doing her a favour, and may very well be improving both of their vacation experiences. This level of nuance is essentially not possible to discover with a vector similarity search.
Our saturation goal therefore became to find the most different examples of nuanced situations such as these. Semantic similarities gave us a recursive process, even within categories of interest breaking them down into large positive examples, negative examples, and special examples. For us, saturation came when further faceting could not meaningfully differentiate groups of documents; as a corollary, we could not find any further examples that could signpost differentiating factors of our themes.
4. Deployment Case Studies